OpenStudy (anonymous):

Mindy and Troy combined ate 9 pieces of the wedding cake. Mindy ate 3 pieces of cake and Troy had 1/4 of the total cake. Write an equation to determine how many pieces of cake (c) there were in total.

OpenStudy (anonymous):

so if mindy ate 3 and they both ate 9 combined, troy ate 6.

OpenStudy (anonymous):

and since troy ate 1/4 of the cake, we know 6=1/4 of the cake

OpenStudy (anonymous):

so how much do you think 4/4 of the cake is? which means the total (c)

OpenStudy (anonymous):

1/4 * 4 =4/4, which is the whole so 6*4=(c)

OpenStudy (anonymous):

okay...

OpenStudy (anonymous):

confused?

OpenStudy (anonymous):

sorta

OpenStudy (anonymous):

how'd you get troy's 6 pieces

OpenStudy (anonymous):

so if they both at 9 combined, and mindy only ate 3, how much did troy eat?

OpenStudy (anonymous):

so if they both ATE** sorry